# QMGeo: Differentially Private Federated Learning via Stochastic Quantization with Mixed Truncated Geometric Distribution

### v1

@@ VII Conclusions
In this paper, we have presented a novel stochastic quantization method $Q_{\text{QMGeo}}(\cdot)$, that utilizes a mixture of truncated geometric distributions to provide randomness for differential privacy. While reducing communication cost through quantization, we also achieve $\epsilon$-DP without the need of any additive noise. In particular, we have provided privacy analysis for $Q_{\text{QMGeo}}(\cdot)$ both in terms of $\epsilon$-DP and RDP, and demonstrated that certain differential privacy levels can be achieved via properly designed stochastic quantization. We have further conducted an optimality gap analysis that mathematically characterizes the convergence performance of the FL framework.

### v2

@@ VII Conclusions
In this paper, we have presented a novel stochastic quantization method $Q_{\text{QMGeo}}(\cdot)$, that utilizes a mixture of truncated geometric distributions to provide randomness for differential privacy. While reducing communication cost through quantization, we also achieve $\epsilon$-DP without the need of any additive noise. In particular, we have provided a privacy analysis for $Q_{\text{QMGeo}}(\cdot)$ both in terms of $\epsilon$-DP and RDP, and demonstrated that certain differential privacy levels can be achieved via properly designed stochastic quantization. We have further conducted an optimality gap analysis that mathematically characterizes the convergence performance of the FL framework.